# Length-Controlled Margin-Based Preference Optimization without Reference Model

### v1

@@ 4.2 Main Results
LMPO achieves competitive performance compared to existing preference optimization methods with controlled length. As shown in Table 1, while all preference optimization algorithms improve over the SFT baseline, LMPO achieves competitive performance compared to existing methods specifically on AlpacaEval 2 and Arena-Hard with controlled length.
AlpacaEval 2: The prompt lengths of LMPO are significantly shorter than those of SimPO in three of the evaluated settings. Notably, in the case of Mistral-Base (7B), LMPO outperforms SimPO by 3.2% in the LC metric, despite utilizing markedly shorter prompt lengths. These results suggest that while LMPO may not lead in terms of LC and WR, its capacity to achieve competitive performance with more efficient prompt lengths positions it as a well-rounded model. It strikes a favorable balance between performance and efficiency, making it particularly suitable for practical applications where both speed and quality are crucial.
Arena-Hard: LMPO achieves the highest win rate while maintaining a shorter prompt length compared to many competitors, making it the most efficient in terms of both performance and prompt length. Its ability to excel in competitive tasks while preserving prompt efficiency positions it as a top choice for complex environments. It is worth noting that the prompt length in the Llama-3-Base (8B) setting is unusually longer than that of other methods. This may be due to the updated Llama-3 tokenizer occasionally introducing two BOS tokens, which can influence the evaluation results.
Overall, LMPO offers a best-in-class combination of strong performance and prompt efficiency, particularly in Arena-Hard, while remaining highly competitive in AlpacaEval 2. Its ability to balance concise outputs with high-quality performance makes it one of the most practical and effective models across these benchmarks.
@@ 5.1 Reduction of probability decrement
Figure 2: The curves of the chosen (top) and rejected (bottom) log-probabilities during the training process in the Llama-3-Base (8B) setting. The red and green curves represent LMPO and SimPO, respectively.
@@ 6 Conclusion
In this paper, we introduce LMPO, which uses a length-controlled margin-based loss function to mitigate length bias and probability reduction. It features a reference-free loss for memory efficiency and a margin-based term with two normalization methods to balance probability control and model performance. Without requiring a reference model, it remains lightweight while effectively reducing length bias and probability decrement. Despite its simplicity, the method achieves competitive results compared to DPO and its variants across multiple benchmarks, including two open-ended benchmarks: AlpacaEval 2, Arena-Hard v0.1 and six conditional benchmarks used in Huggingface open leaderboard V1.
@@
Appendix A Evaluation Details
@@
  • AI2 Reasoning Challenge: A benchmark for evaluating AI scientific reasoning, consisting of 2,590 multiple-choice questions (Clark et al., 2018). Each question tests science knowledge and reasoning, with highly challenging distractors designed to confuse non-expert models.
  • HellaSwag: A benchmark for testing AI commonsense reasoning, consisting of 70,000 multiple-choice questions (Zellers et al., 2019). Each question has a context and four endings, with one correct answer. Adversarial distractors make it highly challenging.
  • MMLU: A benchmark for evaluating AI across several diverse tasks, including reasoning, knowledge, and language understanding (Hendrycks et al., 2020). It consists of over 12,000 multiple-choice questions, testing models’ performance on tasks ranging from general knowledge to specialized domains.
@@
  • TruthfulQA: a benchmark for evaluating AI’s ability to generate truthful and factual answers, consisting of 818 multiple-choice questions (Lin et al., 2021). It tests models’ capacity to provide accurate information across various domains, with distractors designed to confuse models into providing false answers.
@@
  • Winogrande: A benchmark for evaluating AI commonsense reasoning, consisting of 44,000 sentence-pair questions (Sakaguchi et al., 2021). Each question requires selecting the correct word to resolve an ambiguity, with challenging distractors that test subtle reasoning abilities.
  • GSM8K: A benchmark for evaluating AI’s performance on arithmetic problem solving, consisting of 8,000 high-school-level math word problems (Cobbe et al., 2021). It tests models’ ability to reason through multi-step calculations and select the correct solution from multiple choices.
@@
  • AlpacaEval2: An open-ended, AI-driven generation benchmark designed to compare model performance (Li et al., 2023). The dataset comprises 805 diverse questions and evaluates model responses against GPT-4, with GPT-4 serving as the judge (Achiam et al., 2023). Additionally, we include a length-debiased win rate to minimize potential biases favoring longer responses (Dubois et al., 2024).
@@
  • Arena-Hard v0.1: Arena-Hard is an enhanced version of MT-Bench, consisting of 500 high-quality prompts sourced from real user queries (Li et al., 2024). GPT-4(0613) is used as the baseline model, while GPT-4-Turbo serves as the evaluator. We measure the win rate against the baseline model.
@@
We categorize the first six datasets as conditional benchmarks, and the last two as open-ended benchmarks. Conditional benchmarks require the model to produce answers in a specific format, enabling the calculation of exact match scores or accuracy. Open-ended benchmarks, on the other hand, allow for free-form responses, providing more flexibility in evaluating the model’s performance.
@@
For all conditional benchmarks, we employ the well-established evaluation tool lm-evaluation-harness (Gao et al., 2021).And in order to follow Huggingface open leaderboard V1, we use the same version of lm-eval repository. ***lm-eval repository of Huggingface open leaderboard V1: https://github.com/EleutherAI/lm-evaluation-harness/tree/b281b0921b636bc36ad05c0b0b0763bd6dd43463
@@
Table 4: Downstream task evaluation results of tasks on the Huggingface open leaderboard V1.
MMLU (5) ARC (25) HellaSwag (10) TruthfulQA (0) Winograd (5) GSM8K (5) Average
@@
SFT 60.10 58.28 80.76 40.35 76.40 28.13 57.34
DPO 58.48 61.26 83.59 53.06 76.80 21.76 59.16
IPO 60.23 60.84 83.30 45.44 77.58 27.14 59.09
CPO 59.39 57.00 80.75 47.07 76.48 33.06 58.96
KTO 60.90 62.37 84.88 56.60 77.27 38.51 63.42
ORPO 63.20 61.01 84.09 47.91 78.61 42.15 62.83
R-DPO 59.58 61.35 84.29 46.12 76.56 18.12 57.67
SimPO 59.30 61.86 83.42 46.48 77.19 20.92 58.20
LMPO 58.48 61.43 83.61 50.67 76.87 21.91 58.83
@@
SFT 60.40 63.57 84.79 66.81 76.64 40.49 65.45
DPO 60.53 65.36 85.86 66.71 76.80 40.33 65.93
IPO 60.20 63.31 84.88 67.36 75.85 39.42 65.17
CPO 60.36 63.23 84.47 67.38 76.80 38.74 65.16
KTO 60.52 65.78 85.49 68.45 75.93 38.82 65.83
ORPO 60.43 61.43 84.32 66.33 76.80 36.85 64.36
R-DPO 60.71 66.30 86.01 68.22 76.72 37.00 65.82
SimPO 59.42 65.53 86.07 70.56 76.01 34.87 65.41
LMPO 59.53 65.27 86.12 70.30 76.16 30.63 64.67
@@
SFT 64.88 60.15 81.37 45.33 75.77 46.32 62.30
DPO 64.31 64.42 83.87 53.48 76.32 38.67 63.51
IPO 64.40 62.88 80.46 54.20 72.22 22.67 59.47
CPO 64.98 61.69 82.03 54.29 76.16 46.93 64.35
KTO 64.42 63.14 83.55 55.76 76.09 38.97 63.65
ORPO 64.44 61.69 82.24 56.11 77.51 50.04 65.34
R-DPO 64.19 64.59 83.90 53.41 75.93 39.27 63.55
SimPO 63.94 65.02 83.09 59.44 77.42 31.54 63.41
LMPO 63.94 64.68 83.03 57.98 77.90 36.01 63.92
@@
SFT 67.06 61.01 78.57 51.66 74.35 68.69 66.89
DPO 66.88 63.99 80.78 59.01 74.66 49.81 65.86
IPO 66.52 61.95 77.90 54.64 73.09 58.23 65.39
CPO 67.05 62.29 78.73 54.01 73.72 67.40 67.20
KTO 66.38 63.57 79.51 58.15 73.40 57.01 66.34
ORPO 66.41 61.01 79.38 54.37 75.77 64.59 66.92
R-DPO 66.74 64.33 80.97 60.32 74.82 43.90 65.18
SimPO 65.72 62.88 78.30 60.74 73.01 50.19 65.14
LMPO 66.08 61.77 76.81 60.06 72.85 43.14 63.45
Appendix B Downstream Result Analysis

### v2

@@ 4.2 Main Results
LMPO achieves a favorable trade-off between performance and prompt efficiency across multiple benchmarks. As shown in Table 1, while all preference optimization methods improve upon the SFT baseline, LMPO demonstrates competitive results, particularly on AlpacaEval 2 and Arena-Hard, with a clear advantage in controlling prompt length.
AlpacaEval 2. LMPO generates significantly shorter prompts than SimPO in three evaluated settings. For example, in the Mistral-Base (7B) setting, LMPO outperforms SimPO by 3.2% on the LC metric despite using much shorter prompts. Although LMPO may not achieve the highest scores on LC and WR, its ability to maintain competitive performance with shorter outputs highlights its efficiency. This indicates that LMPO achieves a meaningful trade-off between performance and prompt length, making it a practical option for scenarios where both generation quality and inference speed are important.
Arena-Hard. LMPO obtains the highest win rate among all compared methods, while still maintaining shorter prompt lengths. This showcases its effectiveness in more challenging settings, where both accuracy and efficiency are critical. Interestingly, in the Llama-3-Base (8B) configuration, LMPO’s prompt length is noticeably longer. This is likely caused by tokenizer-related artifacts (e.g., the presence of multiple BOS tokens), which can affect the computed length but not the model’s core efficiency.
Overall. LMPO achieves strong performance on both AlpacaEval 2 and Arena-Hard, with particularly notable results on the latter benchmark. The difference in improvements across the two datasets may stem from their distinct characteristics—Arena-Hard contains more complex and diverse tasks than AlpacaEval 2. LMPO’s stronger results on Arena-Hard further confirm its suitability for handling difficult problems, demonstrating its advantage in complex real-world scenarios. These results suggest that LMPO is a practical and effective approach that balances concise outputs with solid performance across diverse evaluation settings.
@@ 5.1 Reduction of probability decrement
Figure 2: The curves of the chosen (top) and rejected (bottom) log-probabilities during the training process in the Llama-3-Base (8B) setting. The blue and orange curves represent LMPO and SimPO, respectively.
@@ 6 Conclusion
In this paper, we introduce LMPO, which uses a length-controlled margin-based loss function to mitigate length bias and probability reduction. It features a reference-free loss for memory efficiency and a margin-based term with two normalization methods to balance probability control and model performance. Without requiring a reference model, it remains lightweight while effectively reducing length bias and probability decrement. Despite its simplicity, the method achieves competitive results compared to DPO and its variants across multiple benchmarks, including two open-ended benchmarks: AlpacaEval 2, Arena-Hard v0.1 and ten conditional benchmarks used in Huggingface open leaderboard V1 and V2.
@@
Appendix A Comprehensive Gradient Analysis and Justification of LMPO
We provide a detailed gradient derivation of the LMPO loss to clarify how it improves separation between winning and losing responses, mitigates length bias, and preserves winning response probabilities during training.
1. LMPO Loss Recap.
The LMPO loss is defined as
$\mathcal{L}_{\text{LMPO}}=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}}\left[\log\left(\frac{1}{1+h\cdot\exp\bigl{(}-d(x,y_{w},y_{l})\bigr{)}}\right)\right]$
where
$d(x,y_{w},y_{l})=$ $\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)$ (14)
$-\lambda\,\overline{m}(x,y_{w},y_{l})$
2. Gradient of the Loss w.r.t. $d$.
Denote the sigmoid function as $\sigma(z)=\frac{1}{1+e^{-z}}$. Then,
$\frac{\partial\mathcal{L}_{\text{LMPO}}}{\partial d}={}$ $\frac{\partial}{\partial d}\log(1+h\cdot e^{-d})$ (15)
[…]
For simplicity, assuming $h=1$,
$\frac{\partial\mathcal{L}_{\text{LMPO}}}{\partial d}=-\frac{\exp(-d)}{1+\exp(-d)}=\sigma(-d)-1.$ (16)
3. Gradient of $d$ w.r.t. Model Parameters $\theta$.
Since $d$ depends on $\pi_{\theta}$ through the log-probabilities and the margin term, its gradient is:
$\nabla_{\theta}d=$ $\frac{\beta}{|y_{w}|}\nabla_{\theta}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\nabla_{\theta}\log\pi_{\theta}(y_{l}|x)$ (17)
$-\lambda\,\nabla_{\theta}\overline{m}(x,y_{w},y_{l})$
4. Gradient of Log-Probability Terms.
Recall that
$\log\pi_{\theta}(y|x)=\sum_{t=1}^{|y|}\log\pi_{\theta}(y_{t}|y_{<t},x).$ (18)
Then,
$\nabla_{\theta}\log\pi_{\theta}(y|x)=\sum_{t=1}^{|y|}\nabla_{\theta}\log\pi_{\theta}(y_{t}|y_{<t},x).$ (19)
Dividing by length $|y|$ normalizes this gradient by sequence length, mitigating length bias by ensuring that longer sequences do not dominate gradient magnitudes merely due to token count.
5. Gradient of the Margin Term $\overline{m}$.
The normalized margin is
$\overline{m}=\frac{m-\alpha_{t}}{\beta_{t}},$ (20)
where $\alpha_{t}$, $\beta_{t}$ are EMA estimates updated as
$\alpha_{t+1}$ $=\gamma\alpha_{t}+(1-\gamma)\mu_{\text{batch}},$ (21)
$\beta_{t+1}$ $=\gamma\beta_{t}+(1-\gamma)\sigma_{\text{batch}},$ (22)
with $\mu_{\text{batch}},\sigma_{\text{batch}}$ computed from the current batch margin $m$ values.
Taking gradient w.r.t. $\theta$,
$\nabla_{\theta}\overline{m}=\frac{\nabla_{\theta}m-\nabla_{\theta}\alpha_{t}}{\beta_{t}}-\frac{(m-\alpha_{t})}{\beta_{t}^{2}}\nabla_{\theta}\beta_{t}.$ (23)
Since $\alpha_{t}$ and $\beta_{t}$ are running averages accumulated over batches and do not directly depend on the current model parameters $\theta$ (they depend on past batches), their gradients $\nabla_{\theta}\alpha_{t}$ and $\nabla_{\theta}\beta_{t}$ can be considered negligible within one batch update, i.e.,
$\nabla_{\theta}\alpha_{t}\approx 0,\quad\nabla_{\theta}\beta_{t}\approx 0,$ (24)
which simplifies the gradient to
$\nabla_{\theta}\overline{m}\approx\frac{\nabla_{\theta}m}{\beta_{t}}.$ (25)
6. Gradient of the Margin $m$.
Recall margin function
$m=(1-p_{w})\cdot\left(1-(p_{w}-p_{l})^{5}\right),$ (26)
where $p_{w}=p_{\theta}(y_{w}|x)$ and $p_{l}=p_{\theta}(y_{l}|x)$.
Computing gradients:
$\nabla_{\theta}m=$ $\nabla_{\theta}(1-p_{w})\cdot\left(1-(p_{w}-p_{l})^{5}\right)$ (27)
$+(1-p_{w})\cdot\nabla_{\theta}\left(1-(p_{w}-p_{l})^{5}\right)$
[…]
$-5(1-p_{w})(p_{w}-p_{l})^{4}\cdot\nabla_{\theta}(p_{w}-p_{l})$
Further,
$\nabla_{\theta}(p_{w}-p_{l})=\nabla_{\theta}p_{w}-\nabla_{\theta}p_{l}.$ (28)
Note that
$\nabla_{\theta}p_{y}=\nabla_{\theta}\pi_{\theta}(y|x)=\pi_{\theta}(y|x)\nabla_{\theta}\log\pi_{\theta}(y|x).$ (29)
7. Interpretation of the Gradient Terms.
  • The negative terms with respect to $\nabla_{\theta}p_{w}$ push to increase $p_{w}$ (winning probability), especially when $p_{w}$ is small or close to $p_{l}$.
  • The terms involving $\nabla_{\theta}p_{l}$ push to decrease $p_{l}$ (losing probability).
  • The exponent 5 in $(p_{w}-p_{l})^{5}$ amplifies the penalty when $p_{w}$ is close to $p_{l}$, increasing gradient magnitude for small preference margins, encouraging better separation.
8. Effect of Length Normalization.
Since $d$ scales log probabilities by $\frac{\beta}{|y|}$, the gradient per token is normalized, which reduces bias towards longer sequences. The scaling factor
$\frac{|y_{w}|+|y_{l}|}{2|y|}$
further adjusts for relative length differences, promoting fairness between candidates.
9. Summary of Gradient Behavior.
Overall, the full gradient of the LMPO loss w.r.t. model parameters $\theta$ is:
$\nabla_{\theta}\mathcal{L}_{\mathrm{LMPO}}={}$ $\mathbb{E}_{(x,y_{w},y_{l})}\Biggl{[}(\sigma(-d+\log h)-1)\cdot$ (30)
$\quad\left(\frac{\beta}{|y_{l}|}\nabla_{\theta}\log\pi_{\theta}(y_{l}|x)\right.$
$\quad\left.-\frac{\beta}{|y_{w}|}\nabla_{\theta}\log\pi_{\theta}(y_{w}|x)+\frac{\lambda}{\beta_{t}}\nabla_{\theta}m\right)\Biggr{]}$
When $h=1$,
$\nabla_{\theta}\mathcal{L}_{\mathrm{LMPO}}={}$ $\mathbb{E}_{(x,y_{w},y_{l})}\Biggl{[}(\sigma(-d)-1)\cdot$ (31)
$\quad\left(\frac{\beta}{|y_{l}|}\nabla_{\theta}\log\pi_{\theta}(y_{l}|x)\right.$
$\quad\left.-\frac{\beta}{|y_{w}|}\nabla_{\theta}\log\pi_{\theta}(y_{w}|x)+\frac{\lambda}{\beta_{t}}\nabla_{\theta}m\right)\Biggr{]}$
This gradient enforces:
  • Increasing $\log\pi_{\theta}(y_{w}|x)$ to boost winning response probability.
  • Decreasing $\log\pi_{\theta}(y_{l}|x)$ to suppress losing response probability.
  • Additional margin-driven gradients to sharpen the preference gap, especially when probabilities are close.
  • Length normalization to prevent long sequence bias.
  • EMA-based normalization for stable training dynamics, preventing abrupt changes in margin scaling.
Conclusion:
The detailed gradient analysis confirms that LMPO effectively helps separate winning and losing responses, alleviates length bias by normalizing gradients per token and by dynamic length scaling, and preserves or even increases winning response probabilities by preventing excessive penalization through stable margin normalization. This results in a more robust and stable preference learning framework for large language models.
Appendix B Evaluation Details
@@
  • AI2 Reasoning Challenge: AI2 Reasoning Challenge is a benchmark designed to evaluate scientific reasoning in AI systems, comprising 2,590 multiple-choice questions. It assesses both factual knowledge and logical reasoning, with carefully crafted distractors that aim to mislead non-expert models.
@@
  • TruthfulQA: TruthfulQA is a benchmark for evaluating the ability of models to generate truthful and factually accurate responses, consisting of 818 multiple-choice questions across various domains. Distractors are intentionally designed to prompt incorrect or misleading answers, providing a robust test of truthfulness.
@@
  • Winogrande: Winogrande is a large-scale commonsense reasoning dataset containing 44,000 sentence-pair questions. Each question requires selecting the correct word to resolve an ambiguity, with challenging distractors that test the model’s ability to perform subtle contextual reasoning.
  • GSM8K: GSM8K is a benchmark for arithmetic problem solving, featuring 8,000 high school-level math word problems. It evaluates a model’s capacity to perform multi-step reasoning and arrive at the correct solution among several answer choices.
  • IFEval: IFEval is a benchmark designed to assess a model’s ability to follow explicit instructions, such as including specific keywords or adhering to a given format. It emphasizes formatting fidelity over content quality, enabling the use of strict evaluation metrics.
  • BBH: BBH(Big Bench Hard) is a curated subset of 23 challenging tasks from the BigBench dataset, targeting complex skills such as multi-step arithmetic, algorithmic reasoning, language understanding, and factual knowledge. The tasks are objectively scored, statistically robust, and align well with human judgment, making BBH a reliable measure of model competence.
  • MATH: MATH consists of high-school level competition math problems drawn from various sources. All items are standardized using LaTeX for equations and Asymptote for diagrams. We retain only level-5 problems, which require solutions in a strict, structured format.
  • GPQA: GPQA is a graduate-level question-answering benchmark developed by PhD experts in domains such as biology, chemistry, and physics. The questions are designed to be accessible to experts but difficult for non-specialists. To preserve its integrity, GPQA is gated and does not provide raw text examples, as per the authors’ guidelines.
  • MuSR: MuSR presents multi-step reasoning challenges based on long-form scenarios ( 1,000 words), such as murder mysteries, spatial reasoning, and team optimization tasks. The dataset demands both complex reasoning and long-range context tracking, with most models performing near chance levels.
  • MMLU-Pro: MMLU-PrO is an improved version of the MMLU benchmark, addressing prior issues like noisy samples and declining difficulty. It increases the number of answer choices (from 4 to 10), raises reasoning requirements, and incorporates expert validation, resulting in a higher-quality and more rigorous benchmark.
@@
  • AlpacaEval2: AlpacaEval 2 is an open-ended generation benchmark comprising 805 diverse prompts used to compare model outputs (Li et al., 2023). GPT-4 is employed as the reference judge (Achiam et al., 2023), and a length-debiased win rate is included to account for potential evaluation biases favoring longer responses (Dubois et al., 2024).
@@
  • Arena-Hard v0.1: Arena-Hard v0.1 is an enhanced version of MT-Bench, including 500 high-quality prompts collected from real user queries (Li et al., 2024). GPT-4 (0613) serves as the baseline model, while GPT-4-Turbo acts as the evaluator. Model performance is assessed based on win rate against the baseline.
@@
We categorize the first ten datasets as conditional benchmarks, and the last two as open-ended benchmarks. Conditional benchmarks require the model to produce answers in a specific format, enabling the calculation of exact match scores or accuracy. Open-ended benchmarks, on the other hand, allow for free-form responses, providing more flexibility in evaluating the model’s performance.
@@
For all conditional benchmarks, we employ the well-established evaluation tool lm-evaluation-harness (Gao et al., 2021).And in order to follow Huggingface open leaderboard V1 and V2, we use the same version of lm-eval repository. ***lm-eval repository of Huggingface open leaderboard V1: https://github.com/EleutherAI/lm-evaluation-harness/tree/b281b0921b636bc36ad05c0b0b0763bd6dd43463 lm-eval repository of Huggingface open leaderboard V2: https://github.com/huggingface/lm-evaluation-harness/tree/adding_all_changess
@@
Table 4: Downstream task evaluation results of tasks on the Huggingface open leaderboard V1 and V2.
MMLU-PRO IFEval BBH GPQA MUSR MATH GSM8K ARC TruthfulQA Winograd Avg. Rank
@@
DPO 26.73 10.49 43.27 28.44 43.65 1.36 21.76 61.26 53.06 76.80 4.5
SLiC 26.52 12.45 42.33 27.93 33.74 1.38 33.74 55.38 48.36 77.35 4.8
IPO 25.87 11.52 40.59 28.15 42.15 1.25 27.14 60.84 45.44 77.58 5.2
KTO 27.51 12.03 43.66 29.45 43.17 2.34 38.51 62.37 56.60 77.27 2.2
CPO 27.04 13.32 42.05 28.45 42.15 2.15 33.06 57.00 47.07 76.48 4.3
SimPO 27.13 10.63 42.94 29.03 39.68 2.49 20.92 61.86 46.48 77.19 4.5
LMPO 28.05 12.15 43.72 30.37 40.61 2.87 22.06 61.95 50.67 77.43 2.4
@@
DPO 26.81 22.89 45.46 28.19 46.43 1.89 35.25 66.89 68.40 76.32 3.6
SLiC 25.69 29.53 45.24 27.04 43.90 1.95 39.65 59.90 65.30 76.32 5.2
IPO 25.75 27.85 43.81 26.61 43.55 2.02 39.42 63.31 67.36 75.85 5.7
KTO 27.46 35.42 45.34 28.19 45.77 2.35 38.80 65.72 68.43 75.91 2.8
CPO 26.85 36.81 45.01 28.15 43.28 2.28 38.74 63.23 67.38 76.80 4.1
SimPO 27.10 37.52 45.70 28.04 44.71 2.19 34.87 65.53 68.40 76.01 3.6
LMPO 26.16 35.94 45.84 28.36 44.84 2.49 34.04 65.57 70.56 76.72 2.7
@@
DPO 31.58 33.61 47.80 32.23 40.48 4.53 38.67 64.42 53.48 76.80 5.4
SLiC 31.11 32.37 46.53 33.29 40.55 3.92 48.82 61.43 54.95 77.27 5.2
IPO 30.18 31.52 46.78 32.61 39.58 4.02 22.67 62.88 54.20 72.22 6.8
KTO 31.16 37.10 47.98 33.72 40.21 4.14 38.97 63.14 55.76 76.09 4.0
CPO 30.95 38.57 47.17 33.15 41.59 4.25 46.93 61.69 54.29 76.16 4.2
SimPO 31.61 37.55 48.38 33.22 40.08 4.23 31.54 65.02 59.42 77.42 3.5
LMPO 31.83 36.58 48.51 31.96 40.32 4.98 36.47 65.13 58.04 77.90 3.0
@@
DPO 35.86 44.57 48.31 31.04 39.02 8.23 49.81 63.99 59.01 74.66 3.0
SLiC 33.25 44.01 47.55 30.52 38.10 8.29 66.57 61.26 53.23 76.16 4.0
IPO 32.97 43.27 46.31 30.95 38.58 8.02 58.23 61.95 54.64 73.09 5.8
KTO 35.00 40.12 47.15 29.70 38.10 7.63 57.01 63.57 58.15 73.40 5.6
CPO 34.56 44.08 48.51 30.08 38.81 7.75 67.40 62.29 54.01 73.72 4.7
SimPO 35.09 43.05 48.95 31.29 39.15 8.16 50.19 62.88 60.74 73.01 3.7
LMPO 36.13 45.33 49.64 29.92 39.29 8.26 43.37 61.77 60.06 72.85 4.4
Appendix C Downstream Result Analysis